# Interpretable Evaluation of AI-Generated Content with Language-Grounded Sparse Encoders

### v1

@@ Acknowledgements
This research was supported by the National University of Singapore and benefited from access to computational resources provided by CogAI4S. We acknowledge the potential presence of unintended biases or conflicts of interest, particularly due to the use of LMMs to interpret the visual patterns. We are deeply grateful to the human annotators who contributed their time and expertise to this project, including Yiming Tang, Qiran Zou, Arash Lagzian, Samson Yu, Harshvardhan Saini and Daksh Mor. Their careful assessments played a vital role in the validation of the LanSE framework. We also thank Xuming Ran for his valuable advice and guidance during the initial stages of the project. Finally, we are grateful to Vaishnavh Nagarajan for his insightful feedback and constructive suggestions that helped shape the direction of this work.
@@ A. Qualitative Examples of LanSE Neurons
In this section, we provide qualitative examples of the interpretable visual patterns used in our evaluation framework, LanSE. These visualizations contain individual features that encodes meaningful explanations and the set of images that can activate them, including both photos and generated images.
Figure 12: Feature examples for the human subgroup of prompt match.
Figure 13: Feature examples for the animal subgroup of prompt match.
Figure 14: Feature examples for the object subgroup of prompt match.
Figure 15: Feature examples for the activity subgroup of prompt match.
Figure 16: Feature examples for the environment subgroup of prompt match.
Figure 17: Feature examples for the style subgroup of visual realism.
Figure 18: Feature examples for the artifact subgroup of visual realism.
Figure 19: Feature examples for the distortion subgroup of physical plausibility.
Figure 20: Feature examples for the structure subgroup of physical plausibility.
@@ C. LMM Analysis Prompts
We use the following prompt to summarize the common features among one given neuron and its corresponding subpopulation that can highly activate it.
Feature Summarization Prompt You are an expert in multimodal feature analysis. {(image)(caption)(image)(caption)…} Analyze the commonalities among these samples. Identify: If there exist one common feature that is possessed by all the instances. Summarize and output exactly one feature, for example: ’[Commonality: Animal wildlife in natural habitats]’ and ’[Commonality: Strawberry-based dessert or dish]’. Only answer in general, do not analyze each image one by one, only generate one single concise phrase. Start answer with ’[Commonality:’. End with ’]’
We use the following prompt to assign the prompt match neurons to their corresponding neuron groups, human, animal, object, activity or environment.
Semantics Feature Categorization Prompt You are a helpful assistant categorizing image features into exactly one of the following groups: ”[human, animal, object, activity, environment].” Given the commonality description: commonality, which of the five categories does it best belong to? Please answer with a single word from this list: [human, animal, object, activity, environment]. Wrap your answer in [].
@@
Realism Feature Categorization Prompt Here is a set of image samples along with their captions: {(image)(caption)(image)(caption)…} They are probably unrealistic for certain reason. One type of reason is the style of the images is intended to be unrealistic, like cartoonish or stamp-like, or painting-like. Another type of reason is visual artifact, like a photo of a real scene but is unrealistic because of the violation of conventions or details. Please answer in the following format: [Style, Explanation: {explanation}] or [Artifact, Explanation: {explanation}] Only give me one answer that is common to all the images and can explain all of them, the explanation should be as short as possible, no more than 20 words.
@@
Physics Feature Categorization Prompt Here is a set of image samples along with their captions: {(image)(caption)(image)(caption)…} They all possess one common pattern that make them physically impossible. One type of features is the distortion of the object surface, like distorted faces or blurred surfaces. Another type of features is the structure of the object, like more than six fingers on a hand, one head on a single body or three legs on a single human. What is the common pattern that make them physically impossible? Please answer in the following format: [Distortion, Explanation: {explanation}] or [Structure, Explanation: {explanation}] Only give me one answer that is common to all the images and can explain all of them, the explanation should be as short as possible, no more than 20 words.
@@ D. Method Extention
Certain failure modes in generative models, such as physics violations (e.g., distorted limbs, extra fingers), are both rare and critical, yet difficult to discover through random sampling. To address this, we develop an interactive human-in-the-loop framework that enables targeted pattern discovery at scale while reducing human annotation cost, as shown in Figure 25.
@@
While LanSE neurons are primarily derived from large-scale, sparsity-driven interpretability modules, the discovery of interpretable features depends on the availability of sufficiently frequent patterns in the training data. However, physics-related errors, such as anatomically impossible structures, are often of low frequency and hard to detect without guidance. Exhaustively humanly reviewing 250,000+ images is infeasible. Therefore, we build an iterative system that leverages weak supervision from lightweight classifiers to bootstrap the collection of relevant samples.
@@
This approach enables the discovery of rare but critical patterns in a scalable manner, cutting down the annotation cost by an order of magnitude compared to exhaustive search.
@@
This generates a new set of interpretable neurons, which are curated via the same LMM-based explanation and categorization pipeline. These neurons form the distortion and structure groups within LanSE’s physical plausibility dimension. By grounding these rare pattern detectors in both language and visual representations, LanSE gains the capacity to diagnose physics violations that would otherwise remain underrepresented.
One thing should be mentioned is that this interactive feature discovery pipeline is based on the intuition that task-related patterns are more prevalent in task-related models. In particular, patterns that is related to physics violations are more likely to be discovered in a model that is used to classify images according to whether they have physics violations or not.
@@
E. Metrics Extention
Here are our mathematical definition and calculative derivations of the four defined metrics, prompt match, visual realism, physical plausibility and content diversity.
Prompt Match: This metric captures semantic mismatches between generated images and prompts, such as missing or extraneous content. Grouped into five semantic categories, the metric is computed as the average number of neurons that diverge across modalities. Higher values indicate less semantic alignments. $\mathbf{prompt}$ can be altered with $\mathbf{human}$, $\mathbf{animal}$, $\mathbf{object}$, $\mathbf{activity}$, or $\mathbf{environment}$.
@@ E. Metrics Extention
Physical Plausibility: This metric detects intuitively erroneous patterns that can make an image physically implausible and unwanted, including distortions of the object surfaces or implausible structures like erroneous anatomies. Higher value suggests more errors generated. $\mathbf{phy}$ can be altered with $\mathbf{distortion}$ or $\mathbf{structure}$.
@@
Content Diversity: This metric measures how diverse a generative model’s outputs are. Neurons in the five prompt match groups and the style group are considered as content neurons, and based on them we compute the average divergence of neurons given two image samples. Higher value suggests better model diversity. Expectation is computed on images ($(\mathbf{x_{i}},\mathbf{y_{i}})$) in the distribution with $|\mathrm{N}_{\mathbf{con}}(\mathbf{x_{i}},\mathbf{y_{i}})|\neq 0$. $\mathbf{con}$ can be altered with $\mathbf{human}$, $\mathbf{animal}$, $\mathbf{object}$, $\mathbf{activity}$, $\mathbf{environment}$ or $\mathbf{style}$.
@@ F. Human Evaluation Protocols
In this section, we detail the procedures we conduct for the human evaluations. Our human evaluation protocol for LansE’s feature accuracy is shown in Figure 26. Human annotators are trained and tasked to identify if the image generally matches with the interpretation from LanSE. We release the human evaluation website for transparency: https://512d-223-109-239-18.ngrok-free.app/.
@@ G. The Choice of LanSE Neuron Groups
1. Semantic Alignment: Does the generated image match the prompt? To evaluate this, we need to detect what semantic content appears in images versus what was requested in text. We organize prompt match neurons into five categories that comprehensively cover visual content:
@@
2. Visual Realism: Does the image look realistic? Unrealistic patterns fall into two distinct categories:
@@
3. Physical Plausibility: Does the image contain patterns that make it unusable? Critical errors that violate physical constraints include:
@@ H. Ablation Studies
To understand the influences of the hyperparameters to the LanSE framework, we conduct two ablation studies on the activation threshold, $\tau$. We evaluate how the number of activated interpretable neurons changes with varying thresholds and how the accuracy of signals in each group change with varying thresholds.
Metrics v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and collect all our defined metrics. The results are listed in Figure 28 and Figure 27.
@@
Figure 29: LanSE’s signal accuracies varying according to the threshold choices. Higher thresholds means stricter constraints on the neural activations and therefore higher accuracies for the signals.
@@ I. Limitations
Language-Describable Patterns Only. LanSE inherently evaluates only patterns that can be described in natural language, introducing an anthropocentric bias. Subtle failures or perceptually important patterns that lack linguistic representations may escape detection. While we argue this aligns with practical requirements, as humans primarily care about describable failures, this constraint limits our ability to capture the full spectrum of possible generation errors.
Dependence on Base Encoder Quality. Our framework’s effectiveness is fundamentally limited by the representations learned by the underlying encoders (CLIP in our implementation). Known biases in CLIP, such as texture bias and limited cultural diversity, propagate through to LanSE evaluations. Future work should explore ensemble approaches using diverse encoders to mitigate these inherited biases.
@@
Evaluation Scope and Coverage. Our framework currently addresses four evaluation dimensions, leaving important aspects like aesthetic quality, cultural appropriateness, and implicit social biases unmeasured. Additionally, certain visual patterns, particularly those involving complex spatial relationships or subtle emotional expressions, may be underrepresented in our discovered features, creating evaluation blind spots.
@@
Static Feature Set. Our current implementation uses a fixed set of discovered features, which may become outdated as generative models evolve and develop new failure modes. While our LMM-based curation pipeline provides some flexibility, adapting to entirely new types of generation errors requires retraining the underlying sparse autoencoders.

### v2

@@ Acknowledgements
This research was supported by the National University of Singapore and benefited from access to computational resources provided by CogAI4S. We are deeply grateful to the human annotators who contributed their time and expertise to this project, including Yiming Tang, Qiran Zou, Arash Lagzian, Samson Yu, Ye Zhang and Yingtao Zhu. Their careful assessments played a vital role in the validation of the LanSE method. Thanks to Xuming Ran for his valuable advice and guidance during the early beginning of the project. We are also grateful to Vaishnavh Nagarajan for his insightful feedback and constructive suggestions that helped shape the direction of this work.
@@ A. Qualitative Examples of LanSE Neurons
In this section, we provide qualitative examples of the interpretable visual patterns used in our evaluation method, LanSE. These visualizations contain individual visual patterns that encodes meaningful explanations and the set of images that can activate them, including both photos and generated images.
Figure 14: Visual pattern examples for the human subgroup of semantic meaning.
Figure 15: Visual pattern examples for the animal subgroup of semantic meaning.
Figure 16: Visual pattern examples for the object subgroup of semantic meaning.
Figure 17: Visual pattern examples for the activity subgroup of semantic meaning.
Figure 18: Visual pattern examples for the environment subgroup of semantic meaning.
Figure 19: Visual pattern examples for the style subgroup of visual realism.
Figure 20: Visual pattern examples for the artifact subgroup of visual realism.
Figure 21: Visual pattern examples for the distortion subgroup of physical plausibility.
Figure 22: Visual pattern examples for the structure subgroup of physical plausibility.
Figure 23: Visual pattern examples for the CXR-LanSE built on chest X-ray images.
@@ C. LMM Analysis Prompts
We use the following prompt to summarize the common patterns among one given neuron and its corresponding subpopulation that can highly activate it.
Pattern Summarization Prompt You are an expert in visual pattern analysis. {(image)(caption)(image)(caption)…} Analyze the commonalities among these samples. Identify: If there exist one common visual pattern that is possessed by all the instances. Summarize and output exactly one visual pattern, for example: ’[Commonality: Animal wildlife in natural habitats]’ and ’[Commonality: Strawberry-based dessert or dish]’. Only answer in general, do not analyze each image one by one, only generate one single concise phrase. Start answer with ’[Commonality:’. End with ’]’
We use the following prompt to assign the semantic meaning neurons to their corresponding neuron groups, human, animal, object, activity or environment.
Semantics Pattern Categorization Prompt You are a helpful assistant categorizing visual patterns into exactly one of the following groups: ”[human, animal, object, activity, environment].” Given the commonality description: commonality, which of the five categories does it best belong to? Please answer with a single word from this list: [human, animal, object, activity, environment]. Wrap your answer in [].
@@
Realism Pattern Categorization Prompt Here is a set of image samples along with their captions: {(image)(caption)(image)(caption)…} They are probably unrealistic for certain reason. One type of reason is the style of the images is intended to be unrealistic, like cartoonish or stamp-like, or painting-like. Another type of reason is visual artifact, like a photo of a real scene but is unrealistic because of the violation of conventions or details. Please answer in the following format: [Style, Explanation: {explanation}] or [Artifact, Explanation: {explanation}] Only give me one answer that is common to all the images and can explain all of them, the explanation should be as short as possible, no more than 20 words.
@@
Physics Pattern Categorization Prompt Here is a set of image samples along with their captions: {(image)(caption)(image)(caption)…} They all possess one common visual pattern that make them physically impossible. One type of visual patterns is the distortion of the object surface, like distorted faces or blurred surfaces. Another type of visual patterns is the structure of the object, like more than six fingers on a hand, one head on a single body or three legs on a single human. What is the common visual pattern that make them physically impossible? Please answer in the following format: [Distortion, Explanation: {explanation}] or [Structure, Explanation: {explanation}] Only give me one answer that is common to all the images and can explain all of them, the explanation should be as short as possible, no more than 20 words.
Additionally, for the hallucination detection part to compare LanSE’s capability with LLMs we use the following prompt.
Hallucination Detection Prompt Analyze this image for physical plausibility errors. Look for: - Wrong number of body parts (fingers, limbs, heads) - Impossible anatomical structures - Objects defying physics - Impossible proportions or configurations Does this image contain physical plausibility errors? Answer: YES or NO Explanation:
@@ D. Method Extention
Certain failure modes in generative models, such as physics violations (e.g., distorted limbs, extra fingers), are both rare and critical, yet difficult to discover through random sampling. To address this, we develop an interactive human-in-the-loop framework that enables targeted visual pattern discovery at scale while reducing human annotation cost, as shown in Figure 28.
@@
While LanSE neurons are primarily derived from large-scale, sparsity-driven interpretability modules, the discovery of visual patterns depends on the availability of sufficiently frequent visual patterns in the training data. However, physics-related errors, such as anatomically impossible structures, are often of low frequency and hard to detect without guidance. Exhaustively humanly reviewing 250,000+ images is infeasible. Therefore, we build an iterative system that leverages weak supervision from lightweight classifiers to bootstrap the collection of relevant samples.
@@
This approach enables the discovery of rare but critical visual patterns in a scalable manner, cutting down the annotation cost by an order of magnitude compared to exhaustive search.
@@
This generates a new set of interpretable neurons, which are curated via the same LMM-based explanation and categorization pipeline. These neurons form the distortion and structure groups within LanSE’s physical plausibility dimension. By grounding these rare visual pattern detectors in both language and visual representations, LanSE gains the capacity to diagnose physics violations that would otherwise remain underrepresented.
One thing should be mentioned is that this interactive visual pattern discovery pipeline is based on the intuition that task-related patterns are more prevalent in task-related models. In particular, visual patterns that is related to physics violations are more likely to be discovered in a model that is used to classify images according to whether they have physics violations or not.
@@
E. Mathematical Definitions of the Four Metrics
Here are our mathematical definition and calculative derivations of the four diagnostic metrics, prompt match, visual realism, physical plausibility and content diversity.
Prompt Match: This metric captures semantic mismatches between generated images and prompts, such as missing or extraneous content. Grouped into five semantic categories, the metric is computed as the average number of neurons that diverge across modalities. Higher values indicate mroe prompt mismatches. $\mathbf{prompt}$ can be altered with $\mathbf{human}$, $\mathbf{animal}$, $\mathbf{object}$, $\mathbf{activity}$, or $\mathbf{environment}$.
@@ E. Mathematical Definitions of the Four Metrics
Physical Plausibility: This metric detects intuitively erroneous visual patterns that can make an image physically implausible and unwanted, including distortions of the object surfaces or implausible structures like erroneous anatomies. Higher value suggests more errors generated. $\mathbf{phy}$ can be altered with $\mathbf{distortion}$ or $\mathbf{structure}$.
@@
Content Diversity: This metric measures how diverse a generative model’s outputs are. Neurons in the five semantic meaning groups and the style group are considered as content neurons, and based on them we compute the average divergence of neurons given two image samples. Higher value suggests better model diversity. Expectation is computed on images ($(\mathbf{x_{i}},\mathbf{y_{i}})$) in the distribution with $|\mathrm{N}_{\mathbf{con}}(\mathbf{x_{i}},\mathbf{y_{i}})|\neq 0$. $\mathbf{con}$ can be altered with $\mathbf{human}$, $\mathbf{animal}$, $\mathbf{object}$, $\mathbf{activity}$, $\mathbf{environment}$ or $\mathbf{style}$.
@@ F. Human Evaluation Protocols
In this section, we detail the procedures we conduct for the human evaluations. Our human evaluation protocol for LansE’s visual pattern detection accuracy is shown in Figure 29. Human annotators are trained and tasked to identify if the image generally matches with the interpretation from LanSE. We release the human evaluation website for transparency: https://512d-223-109-239-18.ngrok-free.app/.
@@ G. The Choice of LanSE Neuron Groups
1. Prompt Match: Does the generated image match the prompt? To evaluate this, we need to detect what semantic content appears in images versus what was requested in text. We organize semantic meaning neurons into five categories that comprehensively cover visual content:
@@
2. Visual Realism: Does the image look realistic? Unrealistic visual patterns fall into two distinct categories:
@@
3. Physical Plausibility: Does the image contain visual patterns that make it unusable? Critical errors that violate physical constraints include:
@@ H. Ablation Studies
To understand the influences of the hyperparameters to the LanSE method, we conduct two ablation studies on the activation threshold, $\tau$. We evaluate how the number of activated interpretable neurons changes with varying thresholds and how the accuracy of signals in each group change with varying thresholds.
Metrics v.s. the Activation Threshold $\tau$ We tune the activation threshold used for the LanSE neurons and collect all the diagnostic metrics. The results are listed in Figure 31 and Figure 30.
@@
Figure 32: LanSE’s signal accuracies varying according to the threshold choices. Higher thresholds means stricter constraints on the identification of visual pattern and therefore higher accuracies for the signals.
@@ I. Limitations
Language-Describable Visual Patterns Only. LanSE inherently evaluates only visual patterns that can be described in natural language, introducing an anthropocentric bias. Subtle failures or perceptually important visual patterns that lack linguistic representations may escape detection. While we argue this aligns with practical requirements, as humans primarily care about describable failures, this constraint limits our ability to capture the full spectrum of possible generation errors.
Dependence on Base Encoder Quality. Our method’s effectiveness is fundamentally limited by the representations learned by the underlying encoders (CLIP in our implementation). Known biases in CLIP, such as texture bias and limited cultural diversity, propagate through to LanSE evaluations. Future work should explore ensemble approaches using diverse encoders to mitigate these inherited biases.
@@
Evaluation Scope and Coverage. Our method currently addresses four evaluation dimensions, leaving important aspects like aesthetic quality, cultural appropriateness, and implicit social biases unmeasured. Additionally, certain visual patterns, particularly those involving complex spatial relationships or subtle emotional expressions, may be underrepresented in our discovered visual patterns, creating evaluation blind spots.
@@
Static Visual Pattern Set. Our current implementation uses a fixed set of discovered patterns, which may become outdated as generative models evolve and develop new failure modes. While our LMM-based curation pipeline provides some flexibility, adapting to entirely new types of generation errors requires retraining the underlying sparse autoencoders.